OpenStudy (anonymous):

integral(-8(csc theta)^4)

OpenStudy (anonymous):

\[-8 \int\limits(\csc \theta)^4 d \theta \]

OpenStudy (anonymous):

I think the first step should be to split up (csctheta)^4

OpenStudy (anonymous):

Do you remember think you know what it would be helpful to break it into?

OpenStudy (anonymous):

it would be (csc theta)^2 * (csc theta)^2 I think

OpenStudy (anonymous):

yea that's perfect because we know the integral of csc^2theta = -cot theta

OpenStudy (anonymous):

but it is useless to have two so I would use a trig identity to make one of them into something else. Remeber csc^2theta = 1 + cot^2 theta

OpenStudy (anonymous):

So what does that turn the problem into?

OpenStudy (anonymous):

\[-8 \int\limits(\csc \theta)^2 +(\csc \theta)^2(\cot \theta)^2 d \theta\]

OpenStudy (anonymous):

good you can make them into seperate integrals. csc^2 theta is a straight forward integral while for the other one you can use a substitution.

OpenStudy (anonymous):

Don't forget about the -8

OpenStudy (anonymous):

Could you confirm my answer? I got \[4\sqrt{x^2 -4}+(2/3)\sqrt{x^2-4}\]

OpenStudy (mathmath333):

thats incorrect

OpenStudy (anonymous):

oh well then.

OpenStudy (anonymous):

where did you get sqrts from?

OpenStudy (anonymous):

oh this was actually a trig substitution integration problem so I substituted the stuff back in. The answer with just the trig functions would be \[-8(-\cot \theta -(1/3)(\cot \theta)^3)\]

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

The original problem was actually \[\int\limits (x^3/\sqrt{x^2-4})dx\]

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{i found a formula for this}\hspace{.33em}\\~\\ \int \csc^nx=\dfrac{\cos x\cdot \csc^{n-1} x}{n-1}+ \dfrac{n-2}{n-1}\int \csc^{n-2} x\,dx }\end{align}\)

OpenStudy (anonymous):

That's a not a formula I learned so I don't know if I am supposed to use that.

OpenStudy (mathmath333):

specially its for the power n

OpenStudy (mathmath333):

i see ur original question is diffferent

OpenStudy (anonymous):